[Brahms-l] Shift report 20050205 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 05 2005 - 08:01:04 EST
1. Five hours store since 0020, problems with injecting and ramping after dump. Another store started at 0745.
2. Runs:
12798 - MRS@45deg@A2255 - 145k tr.3, FS@12deg@B430 - 272k tr.6
12799 - MRS@45deg@A2255 - 268k tr.3, FS@12deg@B430 - 385k tr.6
12800 - MRS@45deg@A2255 - 126k tr.3, FS@12deg@B430 - 180k tr.6
12801 - MRS@45deg@A2255 - 116k tr.3, FS D1 843 B   - 104k tr.6
12803 - MRS@45deg@A2255 -  80k tr.3, FS D2 723 B   - 119k tr.6
Runs 12804 and 12805 - pedestals.
12806 - running - FS@12deg@A861.
3. Some problem with connecting to EVB and IOServer at the beginning of the shift.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 5 08:01:41 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 05 2005 - 08:01:58 EST